# Robust Experimental Signatures of Phase Transitions in the Variational Quantum Eigensolver

### v2

@@ Abstract
The Variational Quantum Eigensolver (VQE) is widely considered to be a promising candidate for a quantum-classical algorithm which could achieve near-term quantum advantage. However, current levels of hardware noise can require extensive application of error-mitigation techniques to achieve reliable computations. In this work, we use several IBM devices to explore a finite-size spin model with multiple ‘phase-like’ regions characterized by distinct ground-state configurations. Using pre-optimized VQE solutions, we demonstrate that in contrast to calculating the energy, where zero-noise extrapolation is required in order to obtain qualitatively accurate yet still unreliable results, calculations of the energy derivative, two-site spin correlation functions, and the fidelity susceptibility yield accurate behavior across multiple regions, even with minimal or no application of error-mitigation approaches. Taken together, these sets of observables could be used to identify level crossings in VQE solutions in a simple and noise-robust manner, with potential near-term application to identifying quantum phase transitions, avoided crossings and non-adiabatic conical intersections in electronic structure calculations.
@@ I Introduction
The promise of quantum computing resides in making classically infeasible computations realizable. The simulation of quantum systems, known to be a classically hard task, yields a particularly natural use-case. However, before the advent of scalable, fault-tolerant quantum computers, extracting relevant results from the currently available ‘noisy intermediate-scale quantum’ (NISQ) devices remains a challenge due to high levels of noise from state preparation, measurement, control, cross talk, etc. These lead to errors which can seriously degrade the results of computations using the experimentally measured data, and vary over time in non-trivial ways [1, 2, 3, 4, 5]. Despite these issues, assertions of the advantage of quantum computation on NISQ devices over classical methods have been made, focused on highly artificial problems that have no immediate real-world application [6, 7, 8, 9]. Furthermore, this preliminary evidence for useful quantum computation before the realization of fully error-corrected quantum devices remains a contentious topic, with many claims being challenged via refinements of existing classical computational schemes [10, 11, 12, 13]. Consequently, a conclusive real-world use case experimentally demonstrating quantum advantage has yet to be identified.
In this pursuit to utilize the capabilities of near- to medium-term available quantum devices for practically significant problems, recent years have seen the development of a large number of techniques to leverage NISQ hardware. Broadly speaking these can be thought of as algorithmic developments and methods of Error Mitigation (EM) which partially compensate for the errors present in computation via pre- and post-processing to bring expectation values of observables closer to their ideal noise-free levels. The topic of EM is a rapidly evolving field: for a thorough recent review of the state of the art, see [14]. On the algorithmic side, one major area of investigation is that of Variational Quantum Algorithms (VQAs), for comprehensive reviews see [15, 16] and references therein.
@@
for a given Hamiltonian $\hat{H}$, where |$\ket{\Psi(p)}=\hat{C}(p)\ket{\Psi_{0}}$ is varied by means of adjusting the parameters $p$ defining the unitary operator $\hat{C}(p)$ acting on the initial state |$\ket{\Psi_{0}}$. This yields an upper bound to the ground-state energy $E_{G}$, attaining the actual value if and only if |$\ket{\Psi(p)}$ is the system’s ground state. The VQE’s capacity to store a low dimensional parameterization of the exponentially large ground state wavefunctions of a system could have a significant impact in one of the most important challenges in condensed matter physics: the characterization of phase diagrams, particularly quantum phase transitions occurring at $T=0$ which are defined by the rearrangement of ground states driven by variations of Hamiltonian parameters. These problems can be extremely challenging for analytical or classical numerical methods to handle due to the quantum critical points (QCPs) defining the transition being characterized by highly entangled states [18], in particular for highly frustrated systems with many competing phases [19].
@@
Subsequently developing methods to identify phases of matter using NISQ hardware and algorithms has received significant attention in recent years. While explorations of equilibrium transitions have seen the application of a plethora of methods including calculating correlation functions in different phases [20, 21, 22], non-local string order parameters and Chern numbers for topological phases [23, 24], tracking changes in fidelity of ground states between Hamiltonian parameters [25], entanglement spectrum [26] or classification by Quantum Convolutional Neural Networks [27, 28, 29], the majority of these studies were either purely classical simulations, digital quantum simulations with ground states prepared by hand from known or classically solved instances, or analogue experiments with limited gate sets. Furthermore those studies which used digitally programmable quantum hardware were generally restricted to a small number of physical qubits, typically less than five.
@@
In contrast, the goal of this paper is to demonstrate that determining phase transitions from VQE solutions on NISQ devices requires the effective utilization of a collection of noise robust observables working in tandem. To accomplish this we apply the VQE to investigate ground-state properties of a frustrated spin model whose ground state manifold is challenging to describe, and utilize real hardware to explore its properties for system sizes up to twelve sites. Utilizing IBM superconducting transmon devices, we particularly focus on the question of which observables can be reliably calculated despite noise in order to determine, potentially without prior knowledge, where transitions occur from these properties alone. However, we find that the noise level of these devices is still prohibitively high when investigating a simple scalar quantity characterizing the system, such as the ground-state energy, failing to even reproduce the qualitative behavior when using parameters pre-optimized via classical simulation. Using EM techniques ameliorates the problem somewhat, but ultimately proves insufficient to produce reliable results or to extract nontrivial features of the system’s phenomenology. Thus, while a narrow focus on the energy to characterize the system seems to show strong limitations for the conclusions available using current hardware, expanding our scope to more noise-robust quantities suggests that this is merely due to an inefficient utilization of data.
@@
Instead, we propose to shift focus to features whose qualitative behavior might be recovered even within NISQ limitations. By exploring several different observables for each VQE solution, we demonstrate that utilizing the measurement data already gathered for the energy calculation in order to construct the first energy derivative and spin-spin correlation functions results in very clean experimental signals of ground state rearrangement under different levels of EM. Furthermore we show that appropriately adapting the Fidelity Susceptibility (FS) to the potentially symmetry breaking expressivity of the VQE ansatz yields a direct measurement of the similarity of different VQE solutions with a high signal to noise ratio, which can be used in the case of topologically differentiated states where local order parameters would fail. We therefore argue that by cross referencing these measurements, one is more capable of accurately inferring whether large fluctuations in a given subset of data are due to corruption from noise or stem from really existing ‘hidden’ noise-free properties of the encoded wavefunction. Consequently, we find that when working within the limitations of present-day quantum devices, focusing on the right kinds of quantities may be just as important as EM. In particular, while noise may wash out even qualitative features of observables such as energy, noise-robust quantities such as the FS, alongside differentiated uses of the data to calculate other observables, may still yield nontrivial information.
@@ II Axial Next-Nearest Neighbor Ising (ANNNI) Model
where $\hat{H}$ is parameterized by $p=(J_{2},B_{x})$. We consider $J_{2}>0$, set $J_{1}=1$ as the unit of energy and use both open and periodic boundary conditions. Combining a nearest-neighbor ferromagnetic Ising coupling with a next-nearest neighbor anti-ferromagnetic coupling, results in a system which displays highly nontrivial features. In particular, varying the strength of the anti-ferromagnetic $J_{2}$ term alongside a transverse field $B_{x}$ results in a $T=0$ phase diagram which exhibits a multitude of quantum phase transitions that pose significant analytical challenges [31]. Although one must of course be in the thermodynamic limit in order to speak of true phases, the structure of these ground states is distinct enough to allow one to identify ‘phase-like’ regions. Thus we use the terms ‘critical point’, ‘phases’ and ‘transitions’ when referencing these different regions of parameter space defined by the structure of the ground-state wavefunction.
@@
The properties of the ANNNI model have been well-studied [30, 31, 19, 32], and in particular its quantum phase structure has been a subject of continuing interest, having been investigated using a variety of techniques, such as quantum convolutional neural networks [33], matrix product states [34], density-matrix renormalization group [35] or perturbative methods [36]. The characteristics of the model have been found to be suitable for describing the properties of the frustrated magnetic compound Ca3Co2O6 [37, 38] which displays characteristics indicative of incommensurate phases formed of periodic order parameters mismatched with respect to the underlying lattice constant [31]. While there are very strong theoretical tools for describing gapped phases in 1D and 2D, if the dominant correlation displays oscillations of this kind, requiring large unit cells to capture, it becomes challenging for 1D methods [39] and virtually impossible for 2D systems, while still being experimentally accessible [40]. Thus given the extensive interest in its properties and the potential for NISQ devices to access highly correlated states, the model forms an ideal testbed for promising approaches to digital quantum simulation, combining a simple system with a well-understood, non-trivial phenomenology.
@@
For $B_{x}=0$, it is trivial to show that there are two ground-state regimes divided by the critical point $J_{2}^{c}=1/2$. For $J_{2}<1/2$ the ground state adopts an energy-degenerate ferromagnetic superposition, namely |$\ket{\Psi}\approx\frac{1}{\sqrt{2}}(\ket{\uparrow}^{\otimes N}+\ket{\downarrow}^{\otimes N})$. Whereas for $J_{2}>1/2$ the next nearest neighbor anti-ferromagnetic term dominates and the ground state adopts what is referred to as a $\braket{2,2}$ configuration, characterized by spin states looking like $|\ket{\uparrow\uparrow\downarrow\downarrow\uparrow\uparrow\downarrow\downarrow\ldots}$. The ground state at the critical point $(J_{2}^{c},0)$ is dominated by fluctuations and is completely degenerate, but numerical studies have shown that lifting the degeneracy by setting $B_{x}\neq 0$, reveals this to be a multi-critical point of a number of phases dependent on boundary conditions and system size $N$, fanning out as $B_{x}$ increases until reaching a paramagnetic phase for $B_{x}\gg J_{2}$.
@@
This rich diversity of phases is characterized by phase-transitions which are generally understood to be second order in nature, with the exception of the phase transition to the $\braket{2,2}$ phase, which is first order [41]. Thus identifying these transitions requires utilizing either an appropriately constructed order parameter, or investigation with correlation functions. As already mentioned, for incommensurate or floating phases this can become a significant hurdle. An alternative metric which has been used to detect ground-state configuration rearrangements in condensed matter and universal properties of critical points in an order-parameter agnostic manner is the Fidelity Susceptibility. In the following section we adopt this measure to partially compensate for the shortcomings of the VQE method.
@@ II.1 Fidelity Susceptibility
Such rotations are not guaranteed to be capable of bringing arbitrary superpositions into perfect alignment with each other. Indeed it is enough to see this by comparing two solutions at the same point. Consider one VQE parameterized solution |$\ket{\Psi_{\text{VQE}}(p)}$ which is in an equal superposition within the degenerate subspace. By construction this state is invariant under all symmetry operations such that $|\langle\Psi_{\text{VQE}}(p)|\hat{U}_{i}|\Psi_{\text{VQE}}(p)\rangle|=1$ for all $\hat{U}_{i}\in\mathcal{U}$. Now consider another solution |$\ket{\tilde{\Psi}_{\text{VQE}}(p)}$ in a completely symmetry broken state, which by construction means that $|\langle\tilde{\Psi}_{\text{VQE}}(p)|\hat{U}_{i}|\tilde{\Psi}_{\text{VQE}}(p)\rangle|<1$ for all $\hat{U}_{i}\neq\hat{I}$. A concrete example for a ferromagnetic ground state with respect to the $\mathbb{Z}_{2}$ symmetry under open boundary conditions would be |$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{\uparrow\uparrow\ldots}+\ket{\downarrow\downarrow\ldots})$, and $|\ket{\tilde{\Psi}}=\ket{\uparrow\uparrow\ldots}$. In this case one could represent the $\mathbb{Z}_{2}$ symmetry as $\hat{U}=X^{\otimes N}$, and $\mathcal{U}=\{I,X^{\otimes N}\}$. It is easy to see that the equal superposition state is, up to a phase, unaffected by the rotations in Eq. (6) and thus $|\braket{\tilde{\Psi}}{\hat{R}(\vec{\theta})}{\tilde{\Psi}}|$ | effectively measures the projection of the $d=2$ dimensional superposition onto one basis state in the symmetry subspace, corresponding to $1/\sqrt{d}$.

### v3

# Noise-Robust Detection of Quantum Phase Transitions

@@ Abstract
Quantum computing allows for the manipulation of highly correlated states whose properties quickly go beyond the capacity of any classical method to calculate. Thus one natural problem which could lend itself to quantum advantage is the study of ground-states of condensed matter models, and the transitions between them. However, current levels of hardware noise can require extensive application of error-mitigation techniques to achieve reliable computations. In this work, we use several IBM devices to explore a finite-size spin model with multiple ‘phase-like’ regions characterized by distinct ground-state configurations. Using pre-optimized Variational Quantum Eigensolver (VQE) solutions, we demonstrate that in contrast to calculating the energy, where zero-noise extrapolation is required in order to obtain qualitatively accurate yet still unreliable results, calculations of the energy derivative, two-site spin correlation functions, and the fidelity susceptibility yield accurate behavior across multiple regions, even with minimal or no application of error-mitigation approaches. Taken together, these sets of observables could be used to identify level crossings in a simple, noise-robust manner which is agnostic to the method of ground state preparation. This work shows promising potential for near-term application to identifying quantum phase transitions, including avoided crossings and non-adiabatic conical intersections in electronic structure calculations.
@@ I Introduction
The promise of quantum computing resides in making classically infeasible computations realizable. The simulation of quantum systems, known to be a classically hard task, yields a particularly natural use-case. However, before the advent of scalable, fault-tolerant quantum computers, extracting relevant results from the currently available ‘noisy intermediate-scale quantum’ (NISQ) devices remains a challenge due to high levels of noise from state preparation, measurement, control, cross talk, etc. These lead to errors which can seriously degrade the results of computations using the experimentally measured data, and vary over time in non-trivial ways [1, 2, 3, 4, 5]. Despite these issues, assertions of the advantage of quantum computation on NISQ devices over classical methods have been made, focused on highly artificial problems that have no immediate real-world application [6, 7, 8, 9]. Furthermore, this preliminary evidence for useful quantum computation before the realization of fully error-corrected quantum devices remains a contentious topic, with many claims being challenged via refinements of existing classical computational schemes [10, 11, 12, 13].
Consequently, a conclusive real-world use case experimentally demonstrating quantum advantage has yet to be identified. Given that in their most naive interpretation, Quantum Computers are devices capable of preparing and manipulating high dimensional states, a natural use case for them is the investigation of the highly entangled and classically intractable ground and dynamical states of correlated condensed matter systems, which can yield novel insights to material structure, function and non-equilibrium properties [14, 15, 16].
In this pursuit to utilize the capabilities of near- to medium-term available quantum devices for such practically significant problems, recent years have seen the development of a large number of techniques to leverage NISQ hardware. Broadly speaking these can be thought of as algorithmic developments and methods of Error Mitigation (EM) which partially compensate for the errors present in computation via pre- and post-processing to bring expectation values of observables closer to their ideal noise-free levels. The topic of EM is a rapidly evolving field: for a thorough recent review of the state of the art, see [17]. On the algorithmic side, one major area of investigation is that of ground state preparation for arbitrary Hamiltonians. Methods such as adiabatic state preparation, quantum phase estimation, Imaginary Time Evolution and the Dissipative Quantum Eigensolver are all active areas of research, for an overview see [18, 19] and references therein.
Variational Quantum Algorithms (VQAs) based on Parameterized Quantum Circuits (PQCs) in particular have garnered a lot of interest on theoretical and pragmatic grounds [20, 21]. On the theoretical side the expressive power of PQCs has been shown to outperform generative neural networks [22], thereby surpassing Neural Quantum States, which themselves can exceed limitations of Tensor Network states widely considered to be a gold standard for area-law entangled states [23]. Pragmatically these algorithms are straightforward to implement on current and near-term devices, and it is primarily for this reason that we utilize this method here.
@@
for a given Hamiltonian $\hat{H}$, where |$\ket{\Psi(p)}=\hat{C}(p)\ket{\Psi_{0}}$ is varied by means of adjusting the parameters $p$ defining the PQC $\hat{C}(p)$ acting on the initial state |$\ket{\Psi_{0}}$. This yields an upper bound to the ground-state energy $E_{G}$, attaining the actual value if and only if |$\ket{\Psi(p)}$ is the system’s ground state.
@@
Regardless of the method by which one prepares the system ground state, the capacity to directly manipulate these states could have a significant impact in one of the most important challenges in condensed matter physics: the characterization of phase diagrams, particularly quantum phase transitions occurring at $T=0$ which are defined by the rearrangement of ground states driven by variations of Hamiltonian parameters. These problems can be extremely challenging for analytical or classical numerical methods to handle due to the quantum critical points (QCPs) defining the transition being characterized by highly entangled states [25], in particular for highly frustrated systems with many competing phases [26].
@@
Subsequently developing methods to identify phases of matter using NISQ hardware and algorithms has received significant attention in recent years. While explorations of equilibrium transitions have seen the application of a plethora of methods including calculating correlation functions in different phases [27, 28, 29], non-local string order parameters and Chern numbers for topological phases [30, 31], tracking changes in fidelity of ground states between Hamiltonian parameters [32], entanglement spectrum [33], coupling systems to external probes which act as witness to transitions [34, 35], machine learning assisted variational optimization [36] or classification by Quantum Convolutional Neural Networks [37, 38, 39], the majority of these studies were either purely classical simulations, digital quantum simulations with ground states prepared by hand from known or classically solved instances, or analogue experiments with limited gate sets. Furthermore those studies which used digitally programmable quantum hardware were generally restricted to a small number of physical qubits, typically less than five, or with limited EM techniques applied.
@@
In contrast, the goal of this paper is to demonstrate that determining phase transitions from states prepared on NISQ devices requires the effective utilization of a collection of noise robust observables working in tandem. To accomplish this we apply the VQE to investigate ground-state properties of a frustrated spin model whose ground state manifold displays a rich variety of highly entangled correlated states, and utilize real hardware to explore its properties for system sizes up to twelve sites. Utilizing IBM superconducting transmon devices, we particularly focus on the question of which observables can be reliably calculated despite noise in order to determine, potentially without prior knowledge, where transitions occur from these properties alone. However, we find that the noise level of these devices is still prohibitively high when investigating a simple scalar quantity characterizing the system, such as the ground-state energy, failing to even reproduce the qualitative behavior when using parameters pre-optimized via classical simulation. Using EM techniques ameliorates the problem somewhat, but ultimately proves insufficient to produce reliable results or to extract nontrivial features of the system’s phenomenology. Thus, while a narrow focus on the energy to characterize the system seems to show strong limitations for the conclusions available using current hardware, expanding our scope to more noise-robust quantities suggests that this is merely due to an inefficient utilization of data.
@@
Instead, we propose to shift focus to features whose qualitative behavior might be recovered even within NISQ limitations. By exploring several different observables for each VQE solution, we demonstrate that utilizing the measurement data already gathered for the energy calculation in order to construct the first energy derivative and spin-spin correlation functions results in very clean experimental signals of ground state rearrangement under different levels of EM. Furthermore we show that appropriately adapting the Fidelity Susceptibility (FS) to the potentially symmetry breaking expressivity of the VQE ansatz yields a direct measurement of the similarity of different VQE solutions with a high signal to noise ratio, which can be used in the case of topologically differentiated states where local order parameters would fail. We therefore argue that by cross referencing these measurements, one is more capable of accurately inferring whether large fluctuations in a given subset of data are due to corruption from noise or stem from really existing ‘hidden’ noise-free properties of the encoded wavefunction. Consequently, we find that when working within the limitations of present-day quantum devices, while noise may wash out even qualitative features of observables such as energy, useful information can still be obtained by focusing on noise-robust quantities such as the FS, alongside other observables, and therefore that the choice of observable may be just as important as EM if one wants to unambiguously demonstrate a meaningful result beyond the limitations of classical computation.
@@ II Axial Next-Nearest Neighbor Ising (ANNNI) Model
where $\hat{H}$ is parameterized by $p=(J_{2},B_{x})$. We consider $J_{2}>0$, set $J_{1}=1$ as the unit of energy and use both open and periodic boundary conditions. Despite its seeming simplicity, the combination of a nearest-neighbor ferromagnetic Ising coupling with a next-nearest neighbor anti-ferromagnetic coupling alongside fluctuations induced by the transverse field $B_{x}$, playing a role analagous to temperature, induces a rich variety of phases, including modulated magnetic orders, commensurate to incommensurate transitions, floating phases, and potential Koster-Thouless infinite order transitions [41, 40, 42, 43, 26, 44, 45]. Although one must of course be in the thermodynamic limit in order to speak of true phases, the structure of these ground states is distinct enough to allow one to identify ‘phase-like’ regions. Thus we use the terms ‘critical point’, ‘phases’ and ‘transitions’ when referencing these different regions of parameter space defined by the structure of the ground-state wavefunction.
Since ANNNI essentially encapsulates the complexity of correlated matter through competition and fluctuation, its phase diagram is a subject of continuing interest in the literature. Yet, there remains no consensus on the number, nature, or location of the phases present [44, 46], despite investigation through a prolific set of methods, including quantum convolutional neural networks [47, 46], matrix product states [48], density-matrix renormalization group (DMRG) [49], quantum monte carlo [42] and perturbative approaches [50]. For example, Beccaria et al. [49] and Nagy [48] both use Matrix Product State based methods and report that all phase transitions converge at one multi-critical point. By contrast the perturbative approach of Chandra and Dasgupta [50] finds that the paramagnetic phase is restricted to sufficiently high $B_{x}$.
@@
One particularly challenging phase to describe is the floating phase, suitable for describing the properties of the frustrated magnetic compound Ca3Co2O6 [51, 52] which displays characteristics indicative of incommensurate phases formed of periodic order parameters mismatched with respect to the underlying lattice constant [41]. While there are very strong theoretical tools for describing gapped phases in 1D and 2D, if the dominant correlation displays oscillations of this kind, requiring large unit cells to capture, it becomes challenging for 1D methods [49, 53] and virtually impossible for 2D systems, while still being experimentally accessible [54]. In the case of the ANNNI model, the study of Beccaria et al. reports that the power law dictating the correlation function which characterizes the phase converges to two different possible exponents depending on the number of sites, and thus even the gold standard of 1D models, DMRG, breaks down. Therefore given the potential for NISQ devices to access highly such correlated states, the model forms an ideal testbed for promising approaches to digital quantum simulation by combining a simple system with a well-understood, non-trivial phenomenology.
@@
The aspects of the phase diagram for this model which are broadly agreed on are that for $B_{x}=0$, it is trivial to show that there are two ground-state regimes divided by the critical point $J_{2}^{c}=1/2$. For $J_{2}<1/2$ the ground state adopts an energy-degenerate ferromagnetic superposition, namely |$\ket{\Psi}\approx\frac{1}{\sqrt{2}}(\ket{\uparrow}^{\otimes N}+\ket{\downarrow}^{\otimes N})$. Whereas for $J_{2}>1/2$ the next nearest neighbor anti-ferromagnetic term dominates and the ground state adopts what is referred to as a $\braket{2,2}$ configuration, characterized by spin states looking like $|\ket{\uparrow\uparrow\downarrow\downarrow\uparrow\uparrow\downarrow\downarrow\ldots}$. The ground state at the critical point $(J_{2}^{c},0)$ is dominated by fluctuations and is completely degenerate, but numerical studies have shown that lifting the degeneracy by setting $B_{x}\neq 0$, reveals this to be a multi-critical point of a number of phases dependent on boundary conditions and system size $N$, fanning out as $B_{x}$ increases until reaching a paramagnetic phase for $B_{x}\gg J_{2}$.
@@
This rich diversity of phases is characterized by phase-transitions which are generally understood to be second order in nature, with the exception perhaps of the phase transition to the $\braket{2,2}$ phase, which could first order [44]. Thus identifying these transitions requires utilizing either an appropriately constructed order parameter, or investigation with correlation functions. As already mentioned, for incommensurate or floating phases this can become a significant hurdle. An alternative metric which has been used to detect ground-state configuration rearrangements in condensed matter and universal properties of critical points in an order-parameter agnostic manner is the Fidelity Susceptibility. Given any circuit which prepares a ground state, one can directly calculate this quantity on a Quantum Computer. However, when dealing the with VQE specifically, we must adapt this method to partially compensate for its shortcomings.
@@ II.1 Fidelity Susceptibility
Such rotations are not guaranteed to be capable of bringing arbitrary superpositions into perfect alignment with each other. Indeed it is enough to see this by comparing two solutions at the same point. Consider one VQE parameterized solution |$\ket{\Psi_{\text{VQE}}(p)}$ which is in an equal superposition within the degenerate subspace. By construction this state is invariant under all symmetry operations such that $|\langle\Psi_{\text{VQE}}(p)|\hat{U}_{i}|\Psi_{\text{VQE}}(p)\rangle|=1$ for all $\hat{U}_{i}\in\mathcal{U}$. Now consider another solution |$\ket{\tilde{\Psi}_{\text{VQE}}(p)}$ in a completely symmetry broken state, which by construction means that $|\langle\tilde{\Psi}_{\text{VQE}}(p)|\hat{U}_{i}|\tilde{\Psi}_{\text{VQE}}(p)\rangle|<1$ for all $\hat{U}_{i}\neq\hat{I}$. A concrete example for a ferromagnetic ground state with respect to the $\mathbb{Z}_{2}$ symmetry under open boundary conditions would be |$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{\uparrow\uparrow\ldots}+\ket{\downarrow\downarrow\ldots})$, and $|\ket{\tilde{\Psi}}=\ket{\uparrow\uparrow\ldots}$. In this case one could represent the $\mathbb{Z}_{2}$ symmetry as $\hat{U}=X^{\otimes N}$, and $\mathcal{U}=\{I,X^{\otimes N}\}$. It is easy to see that the equal superposition state is, up to a phase, unaffected by the rotations in Eq. (6) and thus $|\braket{\tilde{\Psi}}{\hat{R}(\vec{\theta})}{\Psi}|$ | effectively measures the projection of the $d=2$ dimensional superposition onto one basis state in the symmetry subspace, corresponding to $1/\sqrt{d}$.